OpenStudy (anonymous):

Which of the following expressions represents the area of a trapezoid with bases of 3x and 5x and a height of 6x?

OpenStudy (campbell_st):

the area of a trapeziod is A = 1/2 x height x( a+b) a and b are the lengths of the parallel sides... \[A = 1/2\times6x \times(3x + 5x)\] then simplify

OpenStudy (anonymous):

is it 24x2

OpenStudy (campbell_st):

Yes

OpenStudy (anonymous):

thnxx